R. Bachimanchi, K. Baggett, A. Freyberger, R. Geng, C. Hotaver, G. Krafft, A. Kimber, R. Legg, C. Mounts, R. Nelson, T. Plawski, T. Powers, C. Reece, S. Suhring, and H. Wang.
Agenda and Notes
1. Operations Status and Summary/Review (Freyberger)
Some startup problems have occurred. NL20 had enormous heat load even with only single cavity operating. The module will be thermal cycled and is being He processed. NL20 cavity one looks to be trashed. No valve between ion pump and cavity. Suspect pump sputtered when turned on at too high vacuum and junk now in cavity. SRF ops procedure needs to be modified as similar occurrences have occurred in the past.
NL07 also being worked on to recover performance. Looks promising. Some isolated cavities are generating vacuum trips and may need to be de-rated.
Overall goal for near term running is 1040-1060/linac
Many circulator loads failed. Started in injector and then in NLINAC after it started operating. Nothing in the SLINAC yet. Suspect the common LWC system. Could also be old age as the North is older. Troubling rate. Arne pled for focused investigation/answer.
Out of 30 kW klystron spares (for C100 cavities). Two new units on order and 3 will be sent out to be rebuilt.
2. Particulate Movement and Potential Counter (Geng)
3. Plasma Cleaning (Powers)
Leadership of quenches. Offline discussion.
C100 Refurbishment Project for the future?